40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  OMC elog  Not logged in ELOG logo
Entry  Fri Jul 15 22:14:14 2022, Koji, General, General, OMC Lab recovery for the OMC #002 test PXL_20220716_035922673.jpgPXL_20220716_025319391.MP.jpgPXL_20220716_025334324.jpgPXL_20220716_035309066.jpgPXL_20220716_035411435.NIGHT.jpg
    Reply  Tue Jul 26 00:01:59 2022, Koji, General, General, OMC #002 delamination check 2 IMG_1106.JPGIMG_1107.JPGIMG_1110.JPG
Message ID: 414     Entry time: Fri Jul 15 22:14:14 2022     Reply to this: 426
Author: Koji 
Type: General 
Category: General 
Subject: OMC Lab recovery for the OMC #002 test 

- The lab is chilly (18degC)

- Cleaned the lab and the optical table a bit so that the delicate work can be done. The diode test rig (borrowed from Downs - see OMC ELOG 408 and OMC ELOG 409) was removed from the table and brought to the office (to return on Monday)

- The rack electronics were energized.

- The OMC mirrors in use were returned to the cases and stored in the plastic box.

- The optical table was also cleaned. Removed the old Al foils. The table was wiped with IPA

- The OMC #4 was moved to the other part of the table, and then OMC #2 was placed in the nominal place (Attachment 1). Note that the "legs" were migrated from #4 to #2. There are three poles that defines the location of the OMC Transportation

- The lid was removed and the OMC was inspected (Attachment 2). Immediately found some more delamination of the epoxy beneath the cable bracket (Attachment 3). This needs to be taken care of before shipment.

- The cavity was already flashing as usual, and a bit of alignment made the TEM00 flashing.

- The locking was a little tricky because the LB unit seemed to have a gain-dependent offset. After some adjustment, robust locks were achieved. The cavity was then finely adjusted. Attachment 4 shows the CCD image of the reflection. The core of the spot is more or less axisymmetric as usual. There is also a large helo around the spot. I was not aware of this before. I may need to wipe some of the mirrors of the input path.

- As the satisfactory lock was achieved, I called a day by taking a picture of the table (Attachment 5).

Attachment 1: PXL_20220716_035922673.jpg  3.144 MB  Uploaded Fri Jul 15 23:31:29 2022  | Hide | Hide all
PXL_20220716_035922673.jpg
Attachment 2: PXL_20220716_025319391.MP.jpg  2.875 MB  Uploaded Fri Jul 15 23:35:28 2022  | Hide | Hide all
PXL_20220716_025319391.MP.jpg
Attachment 3: PXL_20220716_025334324.jpg  1.947 MB  Uploaded Fri Jul 15 23:36:07 2022  | Hide | Hide all
PXL_20220716_025334324.jpg
Attachment 4: PXL_20220716_035309066.jpg  3.147 MB  Uploaded Fri Jul 15 23:45:01 2022  | Hide | Hide all
PXL_20220716_035309066.jpg
Attachment 5: PXL_20220716_035411435.NIGHT.jpg  2.344 MB  Uploaded Fri Jul 15 23:50:27 2022  | Hide | Hide all
PXL_20220716_035411435.NIGHT.jpg
ELOG V3.1.3-